OpenStudy (anonymous):

are the lines 3x + y = 6 and 2x - 4y = 1 parallel, perpendicular or neither ? also help explaining how to find out .

OpenStudy (anonymous):

If the slopes are equal then parallel If the product of the slopes is -1 then perpendicular

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

a little , can you like show me how to do it ?

OpenStudy (anonymous):

Calculate slope of 3x+4y=6

OpenStudy (anonymous):

Can u?

OpenStudy (anonymous):

thats one of the problems, i dont get it AT ALL D;

OpenStudy (anonymous):

3x+4y=6 SLOPE=-3/4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

-.75 then what?

OpenStudy (anonymous):

Calculate slope of 2x - 4y = 1

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

-.5 ?

OpenStudy (anonymous):

its -2/(-4) =.5

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

okay okay , and they're.. perpendicular then right?

OpenStudy (anonymous):

3x+4y=6 substract -3x from both sides of equal sign -3x+3x+4y=6-3x 4y=-3x+6 divide by 4 4/4 y=-3/4 x+6/4 y= -3/4 x+ 3/2

OpenStudy (anonymous):

Now, neither -.75 is equal to .5 nor -.75*.5 is -1 So, they are neither parallel nor perpendicelar

OpenStudy (anonymous):

... what , okay so explain it to me like you would to a 10 year old . : /

OpenStudy (radar):